OpenStudy (anonymous):

Solve each system by using elimination -x+5y=13 x-y=15 Please show work Thank you

OpenStudy (rizags):

add straight down the rows

OpenStudy (rizags):

you get -x+x+5y-y=13+15

OpenStudy (rizags):

simplifies to 4y=28, or y=7

OpenStudy (rizags):

plug this y back into the original x-y=15, to get x-7=15, or x=22

OpenStudy (rizags):

therefore x=22, and y=7